Northwest Utility Company faces increasing needs for capital. Fortunately, it has an Aa3 credit rating. The corporate tax rate is 30 percent. Northwest’s treasurer is trying to determine the corporation’s current weighted average cost of capital in order to assess the profitability of capital budgeting projects. Historically, the corporation’s earnings and dividends per share have increased about 9.2 percent annually and this should continue in the future. Northwest’s common stock is selling at $65 per share, and the company will pay a $7.50 per share dividend (D1). The company’s $98 preferred stock has been yielding 9 percent in the current market. Flotation costs for the company have been estimated by its investment banker to be $5.00 for preferred stock. The company’s optimum capital structure is 60 percent debt, 10 percent preferred stock, and 30 percent common equity in the form of retained earnings. Refer to the following table on bond issues for comparative yields on bonds of equal risk to Northwest.

Issue Moody’s
Rating
Price Yield to Maturity
Utilities:
Southwest electric power––7 1/4 2023 Aa2 $ 900.18 8.84 %
Pacific bell––7 3/8 2025 Aa3 892.25 8.83
Pennsylvania power & light––8 1/2 2022 A2 975.66 8.88
Industrials:
Johnson & Johnson––6 3/4 2023 Aaa 890.24 8.24 %
Dillard’s Department Stores––7 1/8 2023 A2 970.92 8.44
Marriott Corp.––10 2015 B2 1,040.10 9.66
Answer #1

Weighted average cost of capital = weight of debt*after-tax cost of debt + weight of equity*cost of equity + weight of preferred stock*cost of preferred stock

Weights are given as follows: 60 percent debt, 10 percent preferred stock, and 30 percent common equity in the form of retained earnings.

cost of debt is yield to maturity of the debt or bonds. Northwest Utility Company has an Aa3 credit rating. in the table given in the question on bond issues for comparative yields on bonds of equal risk to Northwest, Pacific bell has an issue under Utilities and its issue has a similar credit rating of Aa3. its yield to maturity is 8.83%.

So, Northwest Utility Company's debt will also have same yield to maturity of 8.83% which will be its cost of debt.

After-tax cost of debt = yield to maturity of the debt/bond*(1-tax rate) = 8.83%*(1-0.30) = 8.83%*0.70 = 6.181%

Cost of Preferred stock is the current market yield on it which is 9% given in the question.

Cost of equity = [expected dividend (D1)/current share price] + dividend growth rate = ($7.50/$65) + 9.2% = 0.1154 + 0.092 = 0.2074 or 20.74%

Weighted average cost of capital = 0.60*6.181% + 0.10*9% + 0.30*20.74% = 3.71% + 0.9% + 6.22% = 10.83%